Project

General

Profile

NDOS Commissioning Meeting 2012/02/13

Attendance: Jon, Mat, Gavin, Andrew, Kurt, Athans, Martin, Peter, Tom, Jarek, Jan, Nick, Susan,

Agenda:

  1. Beam and Detector Update
  2. Installation Plans for the week
  3. DAQ Update and Plans for the week
  4. DCS Update and Plans for the week
  5. AOB

Beam and Detector Report

  • Beam came back last Wednesday morning, but beam spills were not being recorded due to an error in the TDU. Andrew: at end of last week, the TDU became unhappy and had to do a hard reboot.
  • Issue over the weekend: disk space was full, which made RC and possibly other processes very unhappy

Installation plans for this week

  • From Leon: "No firm plans on installation yet. Still looking like some testing for drying this week, and installation of more parts later this month, with a bigger bunch in the second/third week of March."
  • Mat: will be giving tours out at NDOS for EDIT school.

DAQ Update and Plans for next week

  • Jon: RC was updated
  • Peter: cleaned up some sticky tags. Jon: ready to make frozen release? Peter: almost... maybe later this week.
  • Jon: any other tests this week? Kurt/Peter: possibly some DDS tests, but might wait for Far Det. cluster to be rebuilt up north. Kurt: some minor tests, moving of log files, etc.

DCS Update and Plans for next week

  • Athans: nothing to report.

AOB

  • Andrew: new TDUs and DCMs deployment on NDOS won't happen until there is an operational readiness review. Earliest that will be done will be end of the week. 5 DCMs, and 1 TDU master + 1 TDU slave.
  • Jon: open shifts, sign up or you might get assigned a shift!